With a per-worker production function y = k1/2, the steady-state capital stock per worker (k*) as a function of the saving rate (s) is given by:

A) k* = (s/ \delta )2.

B) k* = (\delta/s)2.

C) k* = s/\delta.

D) k* = \delta /s.

Answer is A, I need to understand the process. Thanks
Answer #1

Answer

The correct answer is (A) k* = (s/ \delta )2

Steady state occurs when change in capital per worker(k) = 0 i.e. \Delta k = 0

Change in k(\Delta k = 0) = Investment per worker - depreciation per worker.

At equilibrium C + I = C + S where C = consumption per worker, I = investment per worker and S = saving per worker

At equilibrium C + I = C + S => I = S.

S = saving per worker = sy where s = saving rate and y = output per worker

Hence I = S = sy and as y = k1/2 => I = S = sk1/2

depreciation per worker.= depreciation rate* Initial capital stock = \delta k

Hence Steady state occurs when \Delta k = 0

As discussed above, \Delta k = 0 => Investment per worker - depreciation per worker = 0

=> sk1/2 - \delta k = 0

=> k1/2 = (s/\delta)2

Thus, Steady state level of capital per worker (k*) = (s/\delta)2

Hence, the correct answer is (A) k* = (s/ \delta )2